# Format as a Prior: Quantifying and Analyzing Bias in LLMs for Heterogeneous Data

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly employed in applications that require processing information from heterogeneous formats, including texts, tables, infoboxes, and knowledge graphs. However, systematic biases toward particular formats may undermine LLMs’ ability to integrate heterogeneous data impartially, potentially resulting in reasoning errors and increased risks in downstream tasks. Despite these concerns, it remains uncertain whether such format biases are systematic, which data-level factors contribute to them, and what internal mechanisms in LLMs underlie their emergence.
In this paper, we make the first attempt to investigate and analyze the format bias in LLMs. To systematically investigate the aforementioned questions, we conduct a three-stage empirical study by constructing an heterogeneous data conflict scenario for the exploration of bias. The first stage explores the presence and direction of bias across a diverse range of LLMs. The second stage aims to examine how key data-level factors, including information richness, structure quality, and format type, influence these biases.
The third stage analyzes how format bias emerges within LLMs’ attention patterns and evaluates a lightweight intervention to test its effectiveness. Based on these investigations, we identify three future research directions to reduce format bias: enhancing data pre-processing through format repair and normalization, introducing inference-time interventions such as attention re-weighting, and developing format-balanced training corpora. These directions will support the design of more robust and fair heterogeneous data processing systems.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks (Brown et al. 2020). However, their practical deployment remains constrained by key limitations, including factual inaccuracies (commonly referred to as “hallucinations”) (Ji et al. 2023) and incomplete or outdated knowledge (Petroni et al. 2019). One promising direction to address these issues is to incorporate external knowledge sources into the reasoning process—allowing models to ground their outputs in more accurate, up-to-date, and contextually relevant information (Lewis et al. 2020; Gao et al. 2023; Huang and Chang 2022).
@@
The presence of different formats introduces a critical challenge: LLMs may not treat all formats equally when leveraging these heterogeneous data collaboratively. An LLM with strong format preferences interprets information through a distorted lens, giving undue weight to favored formats regardless of actual relevance. This can affect its ability to reason and synthesize effectively. For example, in clinical decision support, an LLM is provided with both a doctor’s preliminary diagnosis notes and auxiliary examination indicators. As illustrated in Figure 1 (a), when these indicators are presented in table, the LLM (e.g., Qwen3-8b) places disproportionate emphasis on the textual notes while overlooking critical information contained in the table. This results in an incorrect diagnosis. In contrast, as shown in Figure 1 (b), presenting the indicators in a KG enables the LLM to identify abnormalities, such as elevated blood sugar levels, and arrive at the correct diagnosis. Similarly, the aforementioned phenomenon may widely arise in heterogeneous reasoning (Christmann, Saha Roy, and Weikum 2024), which involves knowledge-intensive questions from many domains, such as finance, science, and sports. When key evidence is distributed across texts, tables, infoboxes and KGs, LLMs may focus on information from their preferred formats, potentially overlooking crucial data in others. Such bias can result in incomplete or flawed conclusions and undermine the LLMs’ role as impartial and effective synthesizers of heterogeneous inputs.
Although there have been several studies exploring various types of bias in LLMs, such as bias between multi-modal data (Zhu et al. 2024; Zhang et al. 2025), there is a lack of systematic research on format bias. To address this gap, we present the first comprehensive investigation and analysis of format bias in LLMs. Our study centers on three critical questions: whether such format biases are systematic, which data-level factors contribute to them, and what internal mechanisms in LLMs underlie their emergence.
To systematically investigate the three questions, we conduct a three-stage empirical study by constructing an heterogeneous data conflict scenario for the exploration of bias. The first stage explores the presence and direction of bias across a diverse range of LLMs. The second stage aims to examine how key data-level factors, including information richness, structure quality, and format type, influence these biases. The third stage investigates the emergence of format bias within LLMs’ attention mechanisms and evaluates a lightweight intervention strategy to assess its effectiveness in mitigating such bias.
@@ 2.1 Heterogeneous reasoning
To formalize this challenge, recent benchmarks such as COMPMIX (Christmann, Saha Roy, and Weikum 2024) and CompMix-IR (Min et al. 2024) require cross-format reasoning, stimulating the development of hybrid QA systems that combine structured and unstructured inputs.
Current approaches can be broadly categorized into two types: (1) Unified retrieval frameworks, which abstract away format heterogeneity using shared APIs or embedding spaces (Xia et al. 2025; Min et al. 2024); and (2) LLM-centric pipelines, which enhance downstream reasoning via evidence selection, re-ranking, or modular tool use (Christmann and Weikum 2024; Lehmann et al. 2024; Zhang et al. 2024; Biswal et al. 2024).
@@ 2.2 LLM Behavior under Conflicting Evidence
Recent studies have uncovered a broad range of behavioral biases in LLMs, extending beyond social stereotypes to systematic patterns in reasoning and judgment. A key challenge lies in how LLMs handle conflicts—both between their parametric knowledge (internal beliefs) and in-context evidence, and among competing pieces of evidence (Xu et al. 2024). LLMs tend to favor information aligned with their pre-trained knowledge, even when contradicted by accurate inputs (Jin et al. 2024a; Xie et al. 2023). These LLM biases are shaped by factors such as entity popularity (Xie et al. 2023), event recency (Fang et al. 2023), and evidence frequency (Jin et al. 2024a). Input artifacts also affect LLM behavior, such as preferring self-generated content over retrieved passages (Tan et al. 2024). Related efforts have also examined knowledge conflicts in multi-modal (Zhang et al. 2025; Zhu et al. 2024) and multi-agent settings (Ju et al. 2025), where preference biases and inter-agent inconsistency further complicate reasoning.
Benchmarks like ConflictBank (Su et al. 2024), WikiContradict (Hou et al. 2024), and WhoQA (Pham et al. 2024) have been proposed to evaluate how LLMs handle factual or semantic inconsistencies, especially in ambiguous scenarios. In response, a range of strategies have emerged, including conflict-aware decoding (Yuan et al. 2024; Jin et al. 2024a), counterfactual data augmentation (Fang et al. 2023), internal intervention via attention pruning (Jin et al. 2024b), neuron reweighting (Shi et al. 2024), or prompting LLMs to generate multi-answer responses with source attribution (Shaier, Kobren, and Ogren 2024).
@@ 3 Investigation Framework
This section establishes a framework for analyzing format bias in LLMs, encompassing dataset construction, confounding factor exclusion, and automated response evaluation. The dataset construction process is detailed in Appendix A of the Supplementary Material, with evaluation details in Appendix B of the Supplementary Material.
@@ 3.1 Dataset and Format Construction
We construct our dataset based on ConflictBank (Su et al. 2024), a public corpus designed to evaluate LLM behavior under factual conflicts. We randomly sample 4,000 entries, each containing one factual claim and three counterclaims with corresponding supporting evidence. This results in 12,000 contradiction pairs, as each claim is paired individually with each of the three counterclaims. Each pair consists of two different claims about the same subject and relation, each supported by its own piece of evidence.
@@
Figure 2 provides illustrative examples of these four formats, showing how semantically equivalent information can be presented in structurally distinct ways. These formats reflect the most commonly used representations in prior work on heterogeneous reasoning.
To verify the integrity of format conversions, we conduct manual inspections on a randomly sampled 5% subset of the generated data. We check each sample against two criteria: (1) factual equivalence with the source text and (2) syntactic validity of the target format. The factual accuracy rate was 98.7%, and the syntactic validity rate was 99.3%, reflecting strong data fidelity.
@@ 3.2 Confounding Factor Control
Filtering Internal Knowledge: To ensure that LLM responses are based on external evidence rather than parametric memory, we adopt a filtering procedure consistent with prior work (Gekhman et al. 2024). Each factual claim is tested 16 times by directly querying a given LLM with the corresponding question in a zero-shot setting, and only those samples for which the model fails to reproduce the factual claim in all trials are retained.
Randomizing Evidence Order: To eliminate the known bias introduced by input order (Xie et al. 2023), we randomize the sequence of all evidence segments for each input.
@@ 3.3 Evaluated LLMs
This evaluation covers ten LLMs across six major series: GPT-4o-mini (Achiam et al. 2023), LLaMA-3.1 (8B), Mistral (7B), Qwen3 (8B, 14B, 30B-A3B, 32B) (Team 2025), Gemma-2 (9B, 27B) (Team et al. 2024), and GLM-4 (9B) (GLM et al. 2024). These LLMs span a range of sizes and architectures, enabling cross-family comparison of format-driven biases. To ensure reproducibility and eliminate stochasticity, all evaluations were conducted with a temperature setting of 0.
@@ 3.4 Evaluation Protocol and Metrics
Given the scale of the dataset, we develop an automated evaluation pipeline using an LLM as an adjudicator. We employ GPT-4o-mini as the evaluation LLM. For each response produced by the target LLM, we prompt the judge with explicit evaluation criteria to determine which of the two conflicting claims the response aligns with, designated as Source A and Source B.
Each answer is assessed by GPT-4o-mini across three independent runs to ensure evaluation stability. The final label is determined via majority vote (i.e., at least two of the three judgments must agree). This setup enhances evaluation consistency and guards against variance in LLM judgment. To further validate the reliability of this automated evaluation, we manually annotate a randomly sampled 5% subset of the data and compare the results with the AI-generated labels. We compute the agreement rate between human and LLM judgments, which reached 99.8%, indicating high consistency between human and LLM judgments.
@@ 4 Experimental Results and Analysis
In this section, we present the empirical results of our three-stage investigation. Complete experimental details and raw results are provided in Appendix C of Supplementary Material.
@@ 4.1 Establishing the Existence of Format Bias
The second, and more decisive, pattern is a strong directional bias that emerges when an LLM commits to a single-sided response. Despite considerable variation in architecture and scale, LLMs demonstrate a surprisingly consistent pattern of format preferences. As shown in Figure 4, our cross-model analysis reveals a clear preference hierarchy: semantically rich formats such as texts and KGs are consistently favored over visually structured ones like infoboxes and tables. Furthermore, when we group the data by topical domain, we find that these biases persist across domains, suggesting that the observed biases are robust and generalizable. See Appendix C.7 of the Supplementary Material for detailed domain-level results.
@@ 4.3 Mechanism Behind Format Bias
The results reveal a consistent improvement in the models’ ability to attend to both conflicting sources: DCR significantly increases across all format pairs and all three models (, Wilcoxon signed-rank test; see Figure 7). This suggests that enforcing a more balanced attention distribution during inference effectively encourages the model to integrate information from both input segments, rather than overlooking one entirely.
@@ 5 Discussion
Our results suggest that output-level format preferences in LLMs likely originate from pretraining exposure. Addressing this requires deeper changes at the model level, such as training on format-balanced corpora, incorporating contrastive objectives to penalize format-driven decisions, or adding format-aware processing modules. While these approaches are more costly than inference-time adjustments, they offer a path toward fundamentally reducing format bias.
@@ C.7 Format Preference across Topical Domains
Figures 8 present one heatmap per domain, each summarizing the directional preference across format combinations. The consistent patterns observed across domains suggest that format bias reflects stable inductive biases rather than domain-specific artifacts.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly employed in applications that require processing information from heterogeneous formats, including texts, tables, infoboxes, and knowledge graphs. However, systematic biases toward particular formats may undermine LLMs’ ability to integrate heterogeneous data impartially, potentially resulting in reasoning errors and increased risks in downstream tasks. Yet it remains unclear whether such biases are systematic, which data-level factors drive them, and what internal mechanisms underlie their emergence.
In this paper, we present the first comprehensive study of format bias in LLMs through a three-stage empirical analysis. The first stage explores the presence and direction of bias across a diverse range of LLMs. The second stage examines how key data-level factors influence these biases. The third stage analyzes how format bias emerges within LLMs’ attention patterns and evaluates a lightweight intervention to test its effectiveness. Our results show that format bias is consistent across model families, driven by information richness, structure quality, and representation type, and is closely associated with attention imbalance within the LLMs. Based on these investigations, we identify three future research directions to reduce format bias: enhancing data pre-processing through format repair and normalization, introducing inference-time interventions such as attention re-weighting, and developing format-balanced training corpora. These directions will support the design of more robust and fair heterogeneous data processing systems.
Code — https://github.com/NLPGM/Format-as-a-prior
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks (brown2020language). However, their practical deployment remains constrained by key limitations, including factual inaccuracies (commonly referred to as “hallucinations”) (ji2023survey) and incomplete or outdated knowledge (petroni2019language). One promising direction to address these issues is to incorporate external knowledge sources into the reasoning process—allowing models to ground their outputs in more accurate, up-to-date, and contextually relevant information (lewis2020retrieval; gao2023retrieval; huang2022towards).
@@
The presence of different formats introduces a critical challenge: LLMs may not treat all formats equally when leveraging these heterogeneous data collaboratively. An LLM with strong format preferences interprets information through a distorted lens, giving undue weight to favored formats regardless of actual relevance. This can affect its ability to reason and synthesize effectively.
For example, in clinical decision support, an LLM (e.g., Qwen3-8b) given both textual notes and tabular examination data may overemphasize the text while overlooking key indicators in the table, leading to an incorrect diagnosis. In contrast, as shown in Figure 1, presenting the same information in a knowledge graph enables the model to identify abnormalities and reach the correct conclusion.
When homogeneous inputs are converted into heterogeneous ones with equivalent content, accuracy decreases by 9% and 12% on HotpotQA (yang2018hotpotqa) and MuSiQue (trivedi2022musique) (200 samples each), confirming that format heterogeneity can directly impair reasoning performance. This phenomenon may widely arise in heterogeneous reasoning (christmann2024compmix), where key evidence is distributed across texts, tables, infoboxes, and knowledge graphs. LLMs may focus on information from their preferred formats, potentially overlooking crucial data in others. Such bias can result in incomplete or flawed conclusions and undermine the LLMs’ role as impartial and effective synthesizers of heterogeneous inputs.
Although there have been several studies exploring various types of bias in LLMs, such as bias between multi-modal data (zhu2024unraveling; zhang2025evaluating), there is a lack of systematic research on format bias. To address this gap, we present the first comprehensive investigation and analysis of format bias in LLMs. Our study centers on three critical questions: whether such format biases are systematic, which data-level factors contribute to them, and what internal mechanisms in LLMs underlie their emergence.
To systematically investigate the three questions, we conduct a three-stage empirical study by constructing a heterogeneous data conflict scenario for the exploration of bias. The first stage explores the presence and direction of bias across a diverse range of LLMs. The second stage aims to examine how key data-level factors, including information richness, structure quality, and format type, influence these biases. The third stage investigates the emergence of format bias within LLMs’ attention mechanisms and evaluates a lightweight intervention strategy to assess its effectiveness in mitigating such bias.
Our results reveal that format bias is both systematic and consistent across models, driven primarily by differences in information richness, structure quality, and format type. We further show that such bias originates from imbalanced attention allocation during inference and can be partially mitigated through attention-based interventions.
@@ 2.1 Heterogeneous reasoning
To formalize this challenge, recent benchmarks such as COMPMIX (christmann2024compmix) and CompMix-IR (min2024unihgkr) require cross-format reasoning, stimulating the development of hybrid QA systems that combine structured and unstructured inputs.
Current approaches can be broadly categorized into two types: (1) Unified retrieval frameworks, which abstract away format heterogeneity using shared APIs or embedding spaces (xia2025er; min2024unihgkr); and (2) LLM-centric pipelines, which enhance downstream reasoning via evidence selection, re-ranking, or modular tool use (christmann2024rag; lehmann2024beyond; zhang2024spaghetti; biswal2024text2sql).
@@ 2.2 LLM Behavior under Conflicting Evidence
Recent studies have uncovered a broad range of behavioral biases in LLMs, extending beyond social stereotypes to systematic patterns in reasoning and judgment. A key challenge lies in how LLMs handle conflicts—both between their parametric knowledge (internal beliefs) and in-context evidence, and among competing pieces of evidence (xu2024knowledge). LLMs tend to favor information aligned with their pre-trained knowledge, even when contradicted by accurate inputs (jin2024tug; xie2023adaptive). These LLM biases are shaped by factors such as entity popularity (xie2023adaptive), event recency (fang2023getting), and evidence frequency (jin2024tug). Input artifacts also affect LLM behavior, such as preferring self-generated content over retrieved passages (tan2024blinded). Related efforts have also examined knowledge conflicts in multi-modal (zhang2025evaluating; zhu2024unraveling) and multi-agent settings (ju2025investigating), where preference biases and inter-agent inconsistency further complicate reasoning.
Benchmarks like ConflictBank (su2024conflictbank), WikiContradict (hou2024wikicontradict), and WhoQA (pham2024s) have been proposed to evaluate how LLMs handle factual or semantic inconsistencies, especially in ambiguous scenarios. In response, a range of strategies have emerged, including conflict-aware decoding (yuan2024discerning; jin2024tug), counterfactual data augmentation (fang2023getting), internal intervention via attention pruning (jin2024cutting), neuron reweighting (shi2024ircan), or prompting LLMs to generate multi-answer responses with source attribution (shaier2024adaptive).
@@ 3 Investigation Framework
This section establishes a framework for analyzing format bias in LLMs, encompassing dataset construction, confounding factor exclusion, and automated response evaluation. The dataset construction process is detailed in Appendix A, with evaluation details in Appendix B.
@@ 3.1 Dataset and Format Construction
We construct our dataset based on ConflictBank (su2024conflictbank), a public corpus designed to evaluate LLM behavior under factual conflicts. We randomly sample 4,000 entries, each containing one factual claim and three counterclaims with corresponding supporting evidence. This results in 12,000 contradiction pairs, as each claim is paired individually with each of the three counterclaims. Each pair consists of two different claims about the same subject and relation, each supported by its own piece of evidence.
@@
Figure 2 provides illustrative examples of these four formats, showing how semantically equivalent information can be presented in structurally distinct ways. These formats reflect the most commonly used representations in prior work on heterogeneous reasoning. Manual inspection of 5% samples confirms 98.7% factual and 99.3% syntactic accuracy, reflecting strong data fidelity.
@@ 3.2 Confounding Factor Control
Filtering Internal Knowledge: To ensure that LLM responses are based on external evidence rather than parametric memory, we adopt a filtering procedure consistent with prior work (gekhman2024does). Each factual claim is tested 16 times by directly querying a given LLM with the corresponding question in a zero-shot setting, and only those samples for which the model fails to reproduce the factual claim in all trials are retained.
Randomizing Evidence Order: To eliminate the known bias introduced by input order (xie2023adaptive), we randomize the sequence of all evidence segments for each input.
@@ 3.3 Evaluated LLMs
This evaluation covers ten LLMs across six major series: GPT-4o-mini (achiam2023gpt), LLaMA-3.1 (8B), Mistral (7B), Qwen3 (8B, 14B, 30B-A3B, 32B) (qwen3technicalreport), Gemma-2 (9B, 27B) (team2024gemma), and GLM-4 (9B) (glm2024chatglmfamilylargelanguage). These LLMs span a range of sizes and architectures, enabling cross-family comparison of format-driven biases.
To ensure reproducibility and eliminate stochasticity, all evaluations were conducted in deterministic inference mode (temperature = 0, without sampling randomness).
@@ 3.4 Evaluation Protocol and Metrics
Given the dataset’s scale, we adopt an automated evaluation pipeline using LLMs as adjudicators. Specifically, GPT-4o-mini, GLM-4.5-Air (zeng2025glm), and Qwen-plus are employed to judge which of the two conflicting claims (Source A or Source B) each target response supports.
Each model independently evaluates every sample three times for stability, with the final label per model determined by majority vote. We then compute FPR and DCR for each model and report their averages across the three evaluators. This multi-model setup improves consistency and mitigates variance in individual LLM judgments. Manual checks on a randomly sampled 5% subset show 99.8% agreement between human and averaged LLM judgments, confirming high reliability.
@@ 4 Experimental Results and Analysis
In this section, we present the empirical results of our three-stage investigation. Complete experimental details and raw results are provided in Appendix C.
@@ 4.1 Establishing the Existence of Format Bias
The second, and more decisive, pattern is a strong directional bias that emerges when an LLM commits to a single-sided response. Despite considerable variation in architecture and scale, LLMs demonstrate a surprisingly consistent pattern of format preferences. As shown in Figure 4, our cross-model analysis reveals a clear preference hierarchy: semantically rich formats such as texts and KGs are consistently favored over visually structured ones like infoboxes and tables. Furthermore, when we group the data by topical domain, we find that these biases persist across domains, suggesting that the observed biases are robust and generalizable. See Appendix C.7 for detailed domain-level results.
@@ 4.3 Mechanism Behind Format Bias
The results reveal a consistent improvement in the models’ ability to attend to both conflicting sources: DCR significantly increases across all format pairs and all three models (see Figure 7). This suggests that enforcing a more balanced attention distribution during inference effectively encourages the model to integrate information from both input segments, rather than overlooking one entirely. When applied to heterogeneous-input reasoning tasks, the attention-balancing method also improves performance on downstream RAG-style QA datasets, including HotpotQA and MuSiQue, where accuracy increases by 6.5% and 9.5%, respectively.
@@ 5 Discussion
Format preferences likely originate from pretraining data imbalance. Mitigation may involve training on format-balanced corpora, using contrastive objectives, or incorporating format-aware modules. Though costlier than inference-time methods, such approaches offer a more fundamental solution.
@@ 6 Conclusion
Acknowledgments
We sincerely thank the anonymous reviewers for their insightful comments and helpful suggestions, which helped improve this work.
This work was supported by the National Natural Science Foundation of China (NSFC) (Grant No. 62576256) and the Fundamental Research Funds for the Central Universities, China (Grant No. 2042022dx0001).
@@ A.2 Examples of Format Transformation
Following prior work (min2024unihgkr; biswal2024text2sql; christmann2024rag; xia2025er), we directly feed the transformed structured data into the LLMs without additional adapters or format-specific encoders.
@@ C.7 Format Preference across Topical Domains
Figures 8 and 9 present domain-level results for FPR and DCR, respectively. Each FPR heatmap summarizes the directional preference across format combinations, while the corresponding DCR plots capture the extent to which models jointly acknowledge conflicting evidence. The consistent trends across domains indicate that format bias manifests as a stable and systematic inductive pattern rather than a domain-specific artifact.